• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Eric

    Log entry time 08:02:15 on June12,2002

    Entry number 85933

    keyword=Shift summary


    New Crew on shift: Peter Monaghan, Issam Qattan, Eric Voutier
    **************** Guillaume Batigne

    Kinematics: prepare the spectrometers for kinematics Q2_b30. It was finally decided
    to cancel Q2_a35 because of the too high momentum of the right arm.

    0:00 Pulse beam is sent to the Hall
    0:43 CW beam is going to be sent to the Hall for checks
    1:23 Beam is here at 5uA for Harp scan check
    ....... 1st scan gives 27um x 107um beam spot
    ....... End with 50um x 65um, not much better can be done tonight
    2:15 Start pointing measurement
    2:24 Move to the LD2 15 cm target
    2:35 Target in and recovers from the sleeping mode
    ....... MCC have problems to give the raster size we want, it seems that the number
    have changed with respect to Q1: spot++ tells about 2 x 2.5 for 1.5 x 1.8 on machine
    screen.
    3:12 Start data taking
    3:18 Prepare the target for 100 uA
    3:43 Start production run at Q2_b30
    4:30 Continuous beam break at about 10 mn frequency.
    4:40 We got the maximum available current at that moment that is 95uA.
    6:15 Changes made to the physics_kumac (see entry # 85922) for extracting the
    number of physics events; change all Q2 header files with beam energy value in
    MeV instead of GeV as s
    6:43 ROC14 crash
    6:47 Back in business
    6:51 Change kinematics to Q2_b40
    7:11 Start data taking at Q2_b40
    7:42 ROC14 crash
    7:45 Back in business
    8:00 MCC takes beam away for machine development

    Run History
    --------

    Q2_b30

    1448 ...... Pointing measurement with the Multi foil carbon target
    1449 ...... Spot++ check
    1450 ...... Spot++ check
    1451 ...... Spot++ check
    1452 ...... Short run at 40 uA
    1453 ...... Good un (stop at 770 k to check physics and count rates)
    1454 ...... Good run 1.2M
    1455 ...... Good run
    1456 ...... Efficiency run

    Q2_b40

    1458 ...... ROC14 crash
    1459 ...... Good run (258k stop for maintenance)